At the football game, 3 hamburgers and 6 soft drinks cost $21, and 3 hamburgers and 3 soft drinks cost $15. Which two equations can be used to determine the price of a hamburger and the price of a soft drink?
Let x represent the cost of a hamburger and y represent the cost of a soft drink.

Answer by math_helper(2461)

(1)   3x + 6y = 21  
(2)   3x + 3y = 15

This system can be quickly solved for y by, say, subtracting eq(2) from eq(1):

3y = 6
y = 2 —> 3x = 21-12 = 9 —> x = 3

Cost of items:
Hamburg: $3
Soft drink: $2
——

Check:
3*3 + 6*2 = 9+12 = 21 ✔
3*3 + 3*2 = 9+6 = 15 ✔